OpenStudy (anonymous):

How would you solve this equation: |10-3k|=2

OpenStudy (anonymous):

\[\pm (10 - 3k) = 2\] 10 - 3k = 2 3k = 8 k = 8/3 -10 + 3k = 2 3k = 12 k = 4

OpenStudy (anonymous):

k = 8/3 , 4

OpenStudy (lilg132):

|10 - 3k| = 2 or |10 - 3k| = -2 so -3k = 2 - 10 -3k = -8 -k = -8/3 8/3 = k or 10 - 3k = -2 -3k = -2 -10 -3k = -12 -k = -4 4 = k